# Renewal theory for Brownian motion with stochastically gated targets

### v1

@@ 3.2.1 Correlated gates
$\lim_{\epsilon\rightarrow 0}\sum_{b=0,1}\left[{\bf I}-{\bf M}(\epsilon,s)\right]_{ab}^{-1}\widetilde{Q}(|bL-\epsilon|,s)=\widetilde{q}(s)\sum_{b=0,1}{\bm{\Lambda}}(s)_{ab}^{-1}.$ (3.41)
Finally, substitute equation (3.41) back into equation (3.31) and take the limit $s\rightarrow 0$ using
$\lim_{s\rightarrow 0}\widetilde{q}(s)$ $=\frac{L}{2D},$ (3.42a)
$\lim_{s\rightarrow 0}{\bm{\Lambda}}(s)^{-1}$ $=2L\left(\begin{array}[]{cc}A(\alpha)&-B(\alpha)\\ \\ -B(\alpha)&A(\alpha)\end{array}\right)^{-1}=\frac{2L}{A(\alpha)^{2}-B(\alpha)^{2}}\left(\begin{array}[]{cc}A(\alpha)&B(\alpha)\\ \\ B(\alpha)&A(\alpha)\end{array}\right),$ (3.42h)
@@
$A(\alpha)=1+\frac{\sqrt{2\alpha/D}L}{\tanh\sqrt{2\alpha/D}L},\quad B(\alpha)=1+\frac{\sqrt{2\alpha/D}L}{\sinh\sqrt{2\alpha/D}L}.$ (3.43)
@@
$T(x_{0})$ $=\lim_{s\rightarrow 0}\widetilde{Q}(x_{0},s)=\lim_{s\rightarrow 0}\frac{1-\widetilde{J}(L,s|x_{0})+\widetilde{J}(0,s|x_{0})}{s}$
$=-\partial_{s}\widetilde{J}(L,0|x_{0})+\partial_{s}\widetilde{J}(0,0|x_{0})=\frac{x_{0}(L-x_{0})}{2D}.$ (3.44)
@@ 3.2.2 Uncorrelated gates
$\quad+\int_{0}^{t}\rho^{(\epsilon)}_{k\bar{k}|k_{0},-1}(x,t-\tau|\epsilon)\overline{\Pi}_{-1|\bar{k}_{0}}(\tau)J(L,\tau|x_{0})d\tau.$ (3.45)
@@
$p_{k\bar{k}|k_{0}\bar{k}_{0}}(x,t|x_{0})$ $=\Pi_{kk_{0}}(t)\overline{\Pi}_{\bar{k}\bar{k}_{0}}(t)p(x,t|x_{0}),$ (3.46a)
$J_{k_{0}}(0,t|x_{0})$ $=\Pi_{-1,k_{0}}(t)J(0,t|x_{0}),\quad J_{\bar{k}_{0}}(L,t|x_{0})=\overline{\Pi}_{-1,\bar{k}_{0}}(t)J(L,t|x_{0}).$ (3.46b)
Laplace transforming equation (3.45) we have
$\widetilde{\rho}^{(\epsilon)}_{k\bar{k}|k_{0}\bar{k}_{0}}(x,s|x_{0})$ $=\widetilde{p}_{k\bar{k}|k_{0}\bar{k}_{0}}(x,s|x_{0})-\widetilde{\rho}^{(\epsilon)}_{k\bar{k}|-1,\bar{k}_{0}}(x,s|\epsilon)\widetilde{J}_{k_{0}}(0,s|x_{0})$
$\quad+\widetilde{\rho}^{(\epsilon)}_{k\bar{k}|k_{0},-1}(x,s|L-\epsilon)\widetilde{J}_{\bar{k}_{0}}(L,s|x_{0}).$ (3.47)
@@
$\widetilde{\rho}^{(\epsilon)}_{k\bar{k}|-1,\bar{k}_{0}}(x,s|\epsilon)$ $=\widetilde{p}_{k\bar{k}|-1,k_{0}}(x,s|\epsilon)-\widetilde{\rho}^{(\epsilon)}_{k\bar{k}|-1,\bar{k}_{0}}(x,s|\epsilon)\widetilde{J}_{-1}(0,s|\epsilon)$
$\quad+\widetilde{\rho}^{(\epsilon)}_{k\bar{k}|-1,k_{0}}(x,s|L-\epsilon)\widetilde{J}_{\bar{k}_{0}}(L,s|\epsilon),\quad\bar{k}_{0}=\pm 1,$ (3.48a)
$\widetilde{\rho}^{(\epsilon)}_{k\bar{k}|k_{0},-1}(x,s|L-\epsilon)$ $=\widetilde{p}_{k\bar{k}|k_{0},-1}(x,s|L-\epsilon)-\widetilde{\rho}^{(\epsilon)}_{k\bar{k}|k_{0},-1}(x,s|\epsilon)\widetilde{J}_{k_{0}}(0,s|L-\epsilon)$
$\quad+\widetilde{\rho}^{(\epsilon)}_{k\bar{k}|k_{0},-1}(x,s|L-\epsilon)\widetilde{J}_{-1}(L,s|L-\epsilon),\quad k_{0}=\pm 1.$ (3.48b)
There are thus two major differences from the case of uncorrelated gates. First, the matrix ${\bf M}(\epsilon,s)$ appearing in equation (3.39) becomes four-dimensional. Second, equation (2.2) becomes
@@
$\quad+kk_{0}\bar{k}\bar{k}_{0}\widetilde{p}(x,s+2(\bar{\alpha}+\alpha)|x_{0})\bigg).$ (3.49)
@@ 4 Higher-dimensional renewal equation
$\mathbf{y}\rightarrow\mathbf{a}_{\epsilon}(\mathbf{y})=\mathbf{y}-\epsilon\mathbf{n}(\mathbf{y}),$ (4.1)
where $\mathbf{n}(\mathbf{y})$ is the unit normal directed towards the exterior of $\Omega$. (Shifts tangential to the surface do not contribute to the limit $\epsilon\rightarrow 0$ because of the Dirichlet boundary condition.) The higher-dimensional version of the renewal equation (2.17) takes the form
$\rho^{(\epsilon)}_{k|k_{0}}(\mathbf{x},t|\mathbf{x}_{0})$ $=\Pi_{k|k_{0}}(t)p(\mathbf{x},t|\mathbf{x}_{0})$ (4.2)
$+\int_{\partial\Omega}d\mathbf{y}\int_{0}^{t}d\tau\,\rho^{(\epsilon)}_{k|-1}(\mathbf{x},t-\tau|\mathbf{a}_{\epsilon}(\mathbf{y}))\Pi_{-1|k_{0}}(\tau)J(\mathbf{y},\tau|\mathbf{x}_{0}).$
@@
$\frac{\partial p(\mathbf{x},t|\mathbf{x}_{0})}{\partial t}$ $=D{\bm{\nabla}}^{2}p(\mathbf{x},t|\mathbf{x}_{0}),\quad\mathbf{x}\in\Omega,\qquad p(\mathbf{y},t|\mathbf{x}_{0})=0\quad\mathbf{y}\in\partial\Omega.$ (4.3)
@@
$p_{k|k_{0}}(\mathbf{x},t|\mathbf{x}_{0})=\Pi_{kk_{0}}(t)p(\mathbf{x},t|\mathbf{x}_{0}),\quad J_{k_{0}}(\mathbf{y}|\mathbf{x}_{0},t)=\Pi_{-1,k_{0}}(t)J(\mathbf{y}|x_{0},t).$ (4.4)
Laplace transforming the renewal equation (4.2) using the convolution theorem then gives
$\widetilde{\rho}^{(\epsilon)}_{k|k_{0}}(\mathbf{x},s|\mathbf{x}_{0})=\widetilde{p}_{k|k_{0}}(\mathbf{x},s|\mathbf{x}_{0})+\int_{\partial\Omega}\widetilde{\rho}^{(\epsilon)}_{k|-1}(\mathbf{x},s|\mathbf{a}_{\epsilon}(\mathbf{y}))\widetilde{J}_{k_{0}}(\mathbf{y},s|\mathbf{x}_{0})d\mathbf{y},$ (4.5)
@@
$\widetilde{p}_{k|k_{0}}(\mathbf{x},s|\mathbf{x}_{0})=\frac{1}{2}\bigg(\widetilde{G}(\mathbf{x},s|\mathbf{x}_{0})+kk_{0}\widetilde{G}(\mathbf{x},s+2\alpha|\mathbf{x}_{0})\bigg),$ (4.6)
@@
$\widetilde{J}_{k_{0}}(\mathbf{y},s|\mathbf{x}_{0})$ $=\frac{1}{2}\bigg(\widetilde{J}(\mathbf{y},s|\mathbf{x}_{0})-k_{0}\widetilde{J}(\mathbf{y},s+2\alpha|\mathbf{x}_{0})\bigg)$
$=-\frac{D}{2}\bigg({\bm{\nabla}}\widetilde{G}(\mathbf{y},s|\mathbf{x}_{0})-k_{0}{\bm{\nabla}}\widetilde{G}(\mathbf{y},s+2\alpha|\mathbf{x}_{0})\bigg)\cdot\mathbf{n}(\mathbf{y}).$ (4.7)
@@
$D{\bm{\nabla}}^{2}\widetilde{G}(\mathbf{x},s|\mathbf{x}_{0})-s\widetilde{G}(\mathbf{x},s|\mathbf{x}_{0})$ $=-\delta(\mathbf{x}-\mathbf{x}_{0}),\ \mathbf{x}\in\Omega,\quad\widetilde{G}(\mathbf{y},s|\mathbf{x}_{0})=0\quad\mathbf{y}\in\partial\Omega.$ (4.8)
Since the surface $\partial\Omega$ is spatially extended for $d>1$, the corresponding renewal equation (4.5) involves a spatial integration with respect to points $\mathbf{y}\in\partial\Omega$. However, this can be dealt with using spectral methods along analogous lines to Refs. [8, 44, Bressloff23, 45].
The first step is to set $k_{0}=-1$ and $\mathbf{x}_{0}=\mathbf{a}_{\epsilon}(\mathbf{z})$ in equation (4.5 ) with $\mathbf{z}\in\partial\Omega$, which leads to a Fredholm integral equation of the second kind:
$\widetilde{\rho}^{(\epsilon)}_{k|-1}(\mathbf{x},s|\mathbf{a}_{\epsilon}(\mathbf{z}))$ $=\widetilde{p}_{k|-1}(\mathbf{x},s|\mathbf{a}_{\epsilon}(\mathbf{z}))$
$\quad+\int_{\partial\Omega}\widetilde{\rho}^{(\epsilon)}_{k|-1}(\mathbf{x},s|\mathbf{a}_{\epsilon}(\mathbf{y}))\widetilde{J}_{-1}(\mathbf{y},s|\mathbf{a}_{\epsilon}(\mathbf{z}))d\mathbf{y}.$
It is convenient to rewrite equation (4) in a more compact form. Temporarily suppressing the dependence on $k,\mathbf{x},s$, we define
$h^{(\epsilon)}(\mathbf{z})=\widetilde{p}_{k|-1}(\mathbf{x},s|\mathbf{a}_{\epsilon}(\mathbf{z})),\quad u^{(\epsilon)}(\mathbf{z})=\widetilde{\rho}^{(\epsilon)}_{k|-1}(\mathbf{x},s|\mathbf{a}_{\epsilon}(\mathbf{z})),$ (4.10)
@@
${\mathbb{L}}^{(\epsilon)}[u](\mathbf{z})=\int_{\partial\Omega}u(\mathbf{y})\widetilde{J}_{-1}(\mathbf{y},s|\mathbf{a}_{\epsilon}(\mathbf{z}))d\mathbf{y}.$ (4.11)
@@
$u^{(\epsilon)}(\mathbf{z})=h^{(\epsilon)}(\mathbf{z})+{\mathbb{L}}^{(\epsilon)}[u^{(\epsilon)}](\mathbf{z}),$ (4.12)
which is a Fredholm integral equation of the second kind for $L^{2}$- integrable functions on $\partial{\mathcal{M}}$. Since $\partial{\mathcal{M}}$ is a compact domain, it follows that the linear operator ${\mathbb{L}}^{(\epsilon)}$ has a discrete spectrum. Hence, one way to formally solve equation (4.12) is to perform an eigenfunction expansion of the functions $u^{(\epsilon)}(\mathbf{z})$ and $h^{(\epsilon)}(\mathbf{z})$. Here we will use such a spectral decomposition to establish that there is a well-defined solution in the limit $\epsilon\rightarrow 0$.
@@
$\widetilde{J}_{-1}(\mathbf{y},s|\mathbf{a}_{\epsilon}(\mathbf{z}))=\widetilde{J}_{-1}(\mathbf{y},s|\mathbf{z})-\epsilon\left.{\bm{\nabla}}_{\mathbf{a}}\widetilde{J}_{-1}(\mathbf{y},s|\mathbf{a})\right|_{\mathbf{a}=\mathbf{z}}\cdot\mathbf{n}(\mathbf{z})+O(\epsilon^{2}).$ (4.13)
Note that care has to be taken in defining the boundary flux when the initial condition is also on the boundary, see Ref. [Bressloff23] and Appendix A. In particular, if $\mathbf{z}\in\partial\Omega$ then $\widetilde{J}(\mathbf{x},s|\mathbf{z})\equiv D{\bm{\nabla}}\widetilde{G}(\mathbf{x},s|\mathbf{z})\cdot\mathbf{n}(\mathbf{x})=0$ for all $\mathbf{x}\not\in\partial\Omega$ due to the Dirichlet boundary condition $G(\mathbf{x},s|\mathbf{z})=0$. On the other hand, $\widetilde{J}(\mathbf{y},s|\mathbf{z})=\overline{\delta}(\mathbf{y}-\mathbf{z})$ for $\mathbf{y},\mathbf{z}\in\partial\Omega$, where the Dirac delta function is restricted to the surface $\mathcal{U}$}. Combining equations (4.12) and (4.13) then yields
@@
$\hskip 56.9055pt+\frac{\epsilon D}{2}{\bm{\nabla}}_{\mathbf{z}}\bigg({\bm{\nabla}}_{\mathbf{y}}\widetilde{G}(\mathbf{y},s+2\alpha|\mathbf{z})\cdot\mathbf{n}(\mathbf{y})\bigg)\cdot\mathbf{n}(z)+O(\epsilon^{2})\bigg]d\mathbf{y}.$ (4.14)
@@
${\mathbb{L}}_{s,\alpha}[u^{(\epsilon)}]=\epsilon^{-1}h^{(\epsilon)},$ (4.15)
@@
${\mathbb{L}}_{s,\alpha}[u](\mathbf{z})$ $=-\frac{D}{2}\mathbf{n}(\mathbf{z})\cdot{\bm{\nabla}}_{\mathbf{z}}\int_{\partial\Omega}\bigg({\bm{\nabla}}_{\mathbf{y}}\bigg[\widetilde{G}(\mathbf{y},s|\mathbf{z})+\widetilde{G}(\mathbf{y},s+2\alpha|\mathbf{z})\bigg]\cdot\mathbf{n}(\mathbf{y})\bigg)u(\mathbf{y})d\mathbf{y}$ (4.16)
@@
${\mathbb{L}}_{s,\alpha}v_{n}(\mathbf{z},s,\alpha)=\mu_{n}(s,\alpha)v_{n}(\mathbf{z},s,\alpha),\quad\mathbf{z}\in\partial\Omega.$ (4.17)

### v2

@@ 3.2.1 Correlated gates
$\lim_{\epsilon\rightarrow 0}\sum_{b=0,1}\left[{\bf I}-{\bf M}(\epsilon,s)\right]_{ab}^{-1}\widetilde{Q}(|bL-\epsilon|,s)=\widetilde{q}(s)\sum_{b=0,1}{\bm{\Lambda}}(s)_{ab}^{-1}.$ (2.4gsajafgouafao)
Finally, substitute equation (2.4gsajafgouafao) back into equation (LABEL:ren2gateQ) and take the limit $s\rightarrow 0$ using
$\lim_{s\rightarrow 0}\widetilde{q}(s)$ $=\frac{L}{2D},$ (2.4gsajafgouafapa)
$\lim_{s\rightarrow 0}{\bm{\Lambda}}(s)^{-1}$ $=2L\left(\begin{array}[]{cc}A(\alpha)&-B(\alpha)\\ \\ -B(\alpha)&A(\alpha)\end{array}\right)^{-1}=\frac{2L}{A(\alpha)^{2}-B(\alpha)^{2}}\left(\begin{array}[]{cc}A(\alpha)&B(\alpha)\\ \\ B(\alpha)&A(\alpha)\end{array}\right),$ (2.4gsajafgouafaph)
@@
$A(\alpha)=1+\frac{\sqrt{2\alpha/D}L}{\tanh\sqrt{2\alpha/D}L},\quad B(\alpha)=1+\frac{\sqrt{2\alpha/D}L}{\sinh\sqrt{2\alpha/D}L}.$ (2.4gsajafgouafapaq)
@@
$T(x_{0})$ $=\lim_{s\rightarrow 0}\widetilde{Q}(x_{0},s)=\lim_{s\rightarrow 0}\frac{1-\widetilde{J}(L,s|x_{0})+\widetilde{J}(0,s|x_{0})}{s}$ (2.4gsajafgouafapar)
$=-\partial_{s}\widetilde{J}(L,0|x_{0})+\partial_{s}\widetilde{J}(0,0|x_{0})=\frac{x_{0}(L-x_{0})}{2D}.$
@@ 3.2.2 Uncorrelated gates
$\quad+\int_{0}^{t}\rho^{(\epsilon)}_{k\bar{k}|k_{0},-1}(x,t-\tau|\epsilon)\overline{\Pi}_{-1|\bar{k}_{0}}(\tau)J(L,\tau|x_{0})d\tau.$
@@
$p_{k\bar{k}|k_{0}\bar{k}_{0}}(x,t|x_{0})=\Pi_{kk_{0}}(t)\overline{\Pi}_{\bar{k}\bar{k}_{0}}(t)p(x,t|x_{0}),$ (2.4gsajafgouafapata)
$J_{k_{0}}(0,t|x_{0})=\Pi_{-1,k_{0}}(t)J(0,t|x_{0}),\quad J_{\bar{k}_{0}}(L,t|x_{0})=\overline{\Pi}_{-1,\bar{k}_{0}}(t)J(L,t|x_{0}).$ (2.4gsajafgouafapatb)
Laplace transforming equation (2.4gsajafgouafapas) we have
$\widetilde{\rho}^{(\epsilon)}_{k\bar{k}|k_{0}\bar{k}_{0}}(x,s|x_{0})$ $=\widetilde{p}_{k\bar{k}|k_{0}\bar{k}_{0}}(x,s|x_{0})-\widetilde{\rho}^{(\epsilon)}_{k\bar{k}|-1,\bar{k}_{0}}(x,s|\epsilon)\widetilde{J}_{k_{0}}(0,s|x_{0})$ (2.4gsajafgouafapatau)
$\quad+\widetilde{\rho}^{(\epsilon)}_{k\bar{k}|k_{0},-1}(x,s|L-\epsilon)\widetilde{J}_{\bar{k}_{0}}(L,s|x_{0}).$
@@
$\widetilde{\rho}^{(\epsilon)}_{k\bar{k}|-1,\bar{k}_{0}}(x,s|\epsilon)$ $=\widetilde{p}_{k\bar{k}|-1,k_{0}}(x,s|\epsilon)-\widetilde{\rho}^{(\epsilon)}_{k\bar{k}|-1,\bar{k}_{0}}(x,s|\epsilon)\widetilde{J}_{-1}(0,s|\epsilon)$ (2.4gsajafgouafapatava)
$\quad+\widetilde{\rho}^{(\epsilon)}_{k\bar{k}|-1,k_{0}}(x,s|L-\epsilon)\widetilde{J}_{\bar{k}_{0}}(L,s|\epsilon),\quad\bar{k}_{0}=\pm 1,$
$\widetilde{\rho}^{(\epsilon)}_{k\bar{k}|k_{0},-1}(x,s|L-\epsilon)$ $=\widetilde{p}_{k\bar{k}|k_{0},-1}(x,s|L-\epsilon)-\widetilde{\rho}^{(\epsilon)}_{k\bar{k}|k_{0},-1}(x,s|\epsilon)\widetilde{J}_{k_{0}}(0,s|L-\epsilon)$ (2.4gsajafgouafapatavb)
$\quad+\widetilde{\rho}^{(\epsilon)}_{k\bar{k}|k_{0},-1}(x,s|L-\epsilon)\widetilde{J}_{-1}(L,s|L-\epsilon),\quad k_{0}=\pm 1.$
There are thus two major differences from the case of uncorrelated gates. First, the matrix ${\bf M}(\epsilon,s)$ defined in equation (2.4gsajafgouac) becomes four-dimensional. Second, equation (2.4gsx) becomes
@@
$\hskip 85.35826pt+kk_{0}\bar{k}\bar{k}_{0}\widetilde{p}(x,s+2(\bar{\alpha}+\alpha)|x_{0})\bigg).$ (2.4gsajafgouafapatavaw)
@@ 4 Higher-dimensional renewal equation
$\mathbf{y}\rightarrow\mathbf{a}_{\epsilon}(\mathbf{y})=\mathbf{y}-\epsilon\mathbf{n}(\mathbf{y}),$ (2.4gsajafgouafapatava)
where $\mathbf{n}(\mathbf{y})$ is the unit normal directed towards the exterior of $\Omega$. (Shifts tangential to the surface do not contribute to the limit $\epsilon\rightarrow 0$ because of the Dirichlet boundary condition.) The higher-dimensional version of the renewal equation (2.4gr) takes the form
$\rho^{(\epsilon)}_{k|k_{0}}(\mathbf{x},t|\mathbf{x}_{0})$ $=\Pi_{k|k_{0}}(t)p(\mathbf{x},t|\mathbf{x}_{0})$ (2.4gsajafgouafapatavb)
$\quad+\int_{\partial\Omega}d\mathbf{y}\int_{0}^{t}d\tau\,\rho^{(\epsilon)}_{k|-1}(\mathbf{x},t-\tau|\mathbf{a}_{\epsilon}(\mathbf{y}))\Pi_{-1|k_{0}}(\tau)J(\mathbf{y},\tau|\mathbf{x}_{0}).$
@@
$\frac{\partial p(\mathbf{x},t|\mathbf{x}_{0})}{\partial t}$ $=D{\bm{\nabla}}^{2}p(\mathbf{x},t|\mathbf{x}_{0}),\quad\mathbf{x}\in\Omega,\qquad p(\mathbf{y},t|\mathbf{x}_{0})=0\quad\mathbf{y}\in\partial\Omega.$ (2.4gsajafgouafapatavc)
@@
$p_{k|k_{0}}(\mathbf{x},t|\mathbf{x}_{0})=\Pi_{kk_{0}}(t)p(\mathbf{x},t|\mathbf{x}_{0}),\quad J_{k_{0}}(\mathbf{y}|\mathbf{x}_{0},t)=\Pi_{-1,k_{0}}(t)J(\mathbf{y}|x_{0},t).$ (2.4gsajafgouafapatavd)
Laplace transforming the renewal equation (2.4gsajafgouafapatavb) using the convolution theorem then gives
$\widetilde{\rho}^{(\epsilon)}_{k|k_{0}}(\mathbf{x},s|\mathbf{x}_{0})=\widetilde{p}_{k|k_{0}}(\mathbf{x},s|\mathbf{x}_{0})+\int_{\partial\Omega}\widetilde{\rho}^{(\epsilon)}_{k|-1}(\mathbf{x},s|\mathbf{a}_{\epsilon}(\mathbf{y}))\widetilde{J}_{k_{0}}(\mathbf{y},s|\mathbf{x}_{0})d\mathbf{y},$ (2.4gsajafgouafapatave)
@@
$\widetilde{p}_{k|k_{0}}(\mathbf{x},s|\mathbf{x}_{0})=\frac{1}{2}\bigg(\widetilde{G}(\mathbf{x},s|\mathbf{x}_{0})+kk_{0}\widetilde{G}(\mathbf{x},s+2\alpha|\mathbf{x}_{0})\bigg),$ (2.4gsajafgouafapatavf)
@@
$\widetilde{J}_{k_{0}}(\mathbf{y},s|\mathbf{x}_{0})$ $=\frac{1}{2}\bigg(\widetilde{J}(\mathbf{y},s|\mathbf{x}_{0})-k_{0}\widetilde{J}(\mathbf{y},s+2\alpha|\mathbf{x}_{0})\bigg)$ (2.4gsajafgouafapatavg)
$=-\frac{D}{2}\bigg({\bm{\nabla}}\widetilde{G}(\mathbf{y},s|\mathbf{x}_{0})-k_{0}{\bm{\nabla}}\widetilde{G}(\mathbf{y},s+2\alpha|\mathbf{x}_{0})\bigg)\cdot\mathbf{n}(\mathbf{y}).$
@@
$D{\bm{\nabla}}^{2}\widetilde{G}(\mathbf{x},s|\mathbf{x}_{0})-s\widetilde{G}(\mathbf{x},s|\mathbf{x}_{0})$ $=-\delta(\mathbf{x}-\mathbf{x}_{0}),\ \mathbf{x}\in\Omega,\quad\widetilde{G}(\mathbf{y},s|\mathbf{x}_{0})=0\quad\mathbf{y}\in\partial\Omega.$ (2.4gsajafgouafapatavh)
Since the surface $\partial\Omega$ is spatially extended for $d>1$, the corresponding renewal equation (2.4gsajafgouafapatave) involves a spatial integration with respect to points $\mathbf{y}\in\partial\Omega$. However, this can be dealt with using spectral methods.
The first step is to set $k_{0}=-1$ and $\mathbf{x}_{0}=\mathbf{a}_{\epsilon}(\mathbf{z})$ in equation (2.4gsajafgouafapatave) with $\mathbf{z}\in\partial\Omega$, which leads to a Fredholm integral equation of the second kind:
$\widetilde{\rho}^{(\epsilon)}_{k|-1}(\mathbf{x},s|\mathbf{a}_{\epsilon}(\mathbf{z}))=\widetilde{p}_{k|-1}(\mathbf{x},s|\mathbf{a}_{\epsilon}(\mathbf{z}))+\int_{\partial\Omega}\widetilde{\rho}^{(\epsilon)}_{k|-1}(\mathbf{x},s|\mathbf{a}_{\epsilon}(\mathbf{y}))\widetilde{J}_{-1}(\mathbf{y},s|\mathbf{a}_{\epsilon}(\mathbf{z}))d\mathbf{y}.$ (2.4gsajafgouafapatavi)
It is convenient to rewrite equation (2.4gsajafgouafapatavi) in a more compact form. Temporarily suppressing the dependence on $k,\mathbf{x},s$, we define
$h^{(\epsilon)}(\mathbf{z})=\widetilde{p}_{k|-1}(\mathbf{x},s|\mathbf{a}_{\epsilon}(\mathbf{z})),\quad u^{(\epsilon)}(\mathbf{z})=\widetilde{\rho}^{(\epsilon)}_{k|-1}(\mathbf{x},s|\mathbf{a}_{\epsilon}(\mathbf{z})),$ (2.4gsajafgouafapatavj)
@@
${\mathbb{L}}^{(\epsilon)}[u](\mathbf{z})=\int_{\partial\Omega}u(\mathbf{y})\widetilde{J}_{-1}(\mathbf{y},s|\mathbf{a}_{\epsilon}(\mathbf{z}))d\mathbf{y}.$ (2.4gsajafgouafapatavk)
@@
$u^{(\epsilon)}(\mathbf{z})=h^{(\epsilon)}(\mathbf{z})+{\mathbb{L}}^{(\epsilon)}[u^{(\epsilon)}](\mathbf{z}),$ (2.4gsajafgouafapatavl)
which is a Fredholm integral equation of the second kind for $L^{2}$- integrable functions on $\partial{\mathcal{M}}$. Since $\partial{\mathcal{M}}$ is a compact domain, it follows that the linear operator ${\mathbb{L}}^{(\epsilon)}$ has a discrete spectrum. Hence, one way to formally solve equation (2.4gsajafgouafapatavl) is to perform an eigenfunction expansion of the functions $u^{(\epsilon)}(\mathbf{z})$ and $h^{(\epsilon)}(\mathbf{z})$. Here we will use such a spectral decomposition to establish that there is a well-defined solution in the limit $\epsilon\rightarrow 0$.
@@
$\widetilde{J}_{-1}(\mathbf{y},s|\mathbf{a}_{\epsilon}(\mathbf{z}))=\widetilde{J}_{-1}(\mathbf{y},s|\mathbf{z})-\epsilon\left.{\bm{\nabla}}_{\mathbf{a}}\widetilde{J}_{-1}(\mathbf{y},s|\mathbf{a})\right|_{\mathbf{a}=\mathbf{z}}\cdot\mathbf{n}(\mathbf{z})+O(\epsilon^{2}).$ (2.4gsajafgouafapatavm)
Note that care has to be taken in defining the boundary flux when the initial condition is also on the boundary, see Ref. [54] and Appendix A. In particular, if $\mathbf{z}\in\partial\Omega$ then $\widetilde{J}(\mathbf{x},s|\mathbf{z})\equiv D{\bm{\nabla}}\widetilde{G}(\mathbf{x},s|\mathbf{z})\cdot\mathbf{n}(\mathbf{x})=0$ for all $\mathbf{x}\not\in\partial\Omega$ due to the Dirichlet boundary condition $G(\mathbf{x},s|\mathbf{z})=0$. On the other hand, $\widetilde{J}(\mathbf{y},s|\mathbf{z})=\overline{\delta}(\mathbf{y}-\mathbf{z})$ for $\mathbf{y},\mathbf{z}\in\partial\Omega$, where the Dirac delta function is restricted to the surface $\mathcal{U}$}. Combining equations (2.4gsajafgouafapatavl) and (2.4gsajafgouafapatavm) then yields
@@
$\hskip 85.35826pt+\frac{\epsilon D}{2}{\bm{\nabla}}_{\mathbf{z}}\bigg({\bm{\nabla}}_{\mathbf{y}}\widetilde{G}(\mathbf{y},s+2\alpha|\mathbf{z})\cdot\mathbf{n}(\mathbf{y})\bigg)\cdot\mathbf{n}(z)+O(\epsilon^{2})\bigg]d\mathbf{y}.$
@@
${\mathbb{L}}_{s,\alpha}[u^{(\epsilon)}]=\epsilon^{-1}h^{(\epsilon)},$ (2.4gsajafgouafapatavo)
@@
${\mathbb{L}}_{s,\alpha}[u](\mathbf{z})$ $=-\frac{D}{2}\mathbf{n}(\mathbf{z})\cdot{\bm{\nabla}}_{\mathbf{z}}\int_{\partial\Omega}\bigg({\bm{\nabla}}_{\mathbf{y}}\bigg[\widetilde{G}(\mathbf{y},s|\mathbf{z})+\widetilde{G}(\mathbf{y},s+2\alpha|\mathbf{z})\bigg]\cdot\mathbf{n}(\mathbf{y})\bigg)u(\mathbf{y})d\mathbf{y}$
@@
${\mathbb{L}}_{s,\alpha}v_{n}(\mathbf{z},s,\alpha)=\mu_{n}(s,\alpha)v_{n}(\mathbf{z},s,\alpha),\quad\mathbf{z}\in\partial\Omega.$ (2.4gsajafgouafapatavq)